OpenStudy (anonymous):

Emily invested $6500 at 3% interest compounded annually. How much interest will she earn in 5 years? A. $975.00 B. $1035.28 C. $7475.00 D. $7535.28

OpenStudy (anonymous):

6500 is how much times 100?

OpenStudy (anonymous):

Take you original amount. Times 0.03. Add that the original. Then take that new sum time 0.03 and add that number to the answer from the first step. Repeat until you have compounded 5 times

OpenStudy (anonymous):

is that hire purchase or depreciation ?

OpenStudy (jdoe0001):

\(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ P=principal\implies 6,500\\ r=rate\implies 3\%\implies \frac{3}{100}\implies 0.03\\ n=periods\textit{ per year, annually means 1period per year}\\ t=years\implies 5\\ A=P\left(1+\frac{r}{n}\right)^{nt}\implies A=6,500\left(1+\frac{0.03}{1}\right)^{1\cdot 5}\)

OpenStudy (anonymous):

I could not remember the formula for interest! But thank you @jdoe0001 !

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

D. A=P(1+r/100)x number of years

OpenStudy (anonymous):

it is b i swear ijust took the test it is b